# The DECADE cosmic shear project III: validation of analysis pipeline using spatially inhomogeneous data

### v1

@@ 3.3 Shear and redshift calibration uncertainties
The two main sources of systematic uncertainties on our measurements are those associated with the shear estimate and the redshift distribution estimate. We follow the implementation from DES Y3 (Secco & Samuroff et al. 2022; Amon et al. 2022) — which is consistent with those of numerous other works (e.g., Asgari et al. 2021; Dalal et al. 2023) — in the modeling of these uncertainties.
For any shear measurement algorithm, a “calibration” step is needed for the measured shear. This calibration is obtained either from simulations or the data itself, and there is an associated uncertainty with the calibration estimate. In the weak lensing regime, the residual impact of any biases on the measured shear, $\gamma_{\textrm{obs}}$, is modeled with both multiplicative $m_{i}$ and additive components $c_{i}$ (Huterer et al. 2006; Heymans et al. 2006; Bridle & King 2007), that scale the true shear $\gamma_{\textrm{true}}$ as
@@
where $m_{i}$ and $c_{i}$ are constants corresponding to the $i^{\rm th}$ redshift bin. The additive, residual bias often has a sufficiently small amplitude to not be relevant. Furthemore, the shear measurements of our catalog are mean-subtracted (Paper I) following the approach in other works (e.g., Zuntz & Sheldon et al. 2018; Gatti & Sheldon et al. 2021; Yamamoto & Becker et al. 2025). This naturally corrects for such additive biases in the data. The multiplicative residual bias, however, scales the amplitude of $\xi_{\rm\pm}$ and thus propagates directly to key cosmological constraints. As a result, accurate estimates of $m_{i}$ are a necessity. We have made such estimates for our shear catalog using a suite of simulated images of the DECADE survey, as described in detail in Paper I. The simulations inform priors for $m_{i}$ which are listed in Table 1.
For the redshift distribution, we also follow DES Y3 in using a shift in the mean redshift, $\Delta z_{i}$, per bin that then shifts the measured $n(z)$ for each bin $i$ as
$n_{i}(z)=n_{\rm fid,i}\left(z-\Delta z_{i}\right).$ (6)
This parametrization keeps the shape of the $n(z)$ unchanged. While one could explore more complicated parametrizations of the $n(z)$ uncertainties, the above model is sufficient for the statistical power of our data (which is similar to DES Y3). In Paper II, we derived the priors on $\Delta z_{i}$ using the self-organizing map photo-z (SOMPZ) method while accounting for a variety of uncertainty contributions including cosmic variance, shot noise, redshift bias, and photometric zeropoint uncertainties. In the Appendix A of Paper II, we also derived estimates of the $n(z)$ from combining SOMPZ with an independent estimate, clustering redshifts (WZ). The latter method includes information from the spatial clustering between our shear sample and spectroscopic sources. The former, SOMPZ-only method is our fiducial choice, while we use the latter SOMPZ plus WZ method as a cross-check. The resulting priors on $\Delta z_{i}$ from the fiducial setting are listed in Table 1.
@@ 4.1 Scale cuts
As we discussed in Section 3.1, baryons can influence the matter power spectrum on a wide range of scales, presenting a significant source of uncertainty. For example, feedback processes from AGN and supernovae can heat environments and suppress matter clustering, while metal enrichment may offer cooling channels that increase small-scale clustering (e.g., Chisari et al. 2018; Schneider et al. 2019). To reduce the impact of these baryonic imprints, we follow Secco & Samuroff et al. (2022) and Amon et al. (2022) in removing shear measurements whose signal can have a significant contribution from baryon evolution. In particular, we use a simulated data vector derived with the matter-power spectrum measured in the “AGN” model of the OverWhelmingly Large Simulations (OWLS, Schaye et al. 2010, see their Table 2) as a representative example for the baryonic imprints/contamination. The current estimates of the power-spectrum suppression from data are consistent with this OWLS model Bigwood et al. (e.g., 2024, see their Figure 1 and 4).
Our scale cuts are set by requiring the contaminated and uncontamined data vectors to differ by $\Delta\chi^{2}<0.3$. This roughly follows the scale-cuts procedure used in DES Y3. In this work, we derive the scale cuts with the following steps, starting from the unmasked data vector:
@@
  • Of the 20 selected points from the previous step, we find the data point with the highest $\Delta\chi^{2}$ and remove it. We repeat this iteratively until the $\Delta\chi^{2}$ between the contaminated and uncontaminated data vectors, after masking/scale cuts, is $\leq 0.3$.
The $\xi_{-}$ shear measurements are more heavily affected by baryon contamination and thus require more impactful scale cuts. In practice we remove a similar number of points as the fiducial DES Y3 scale cuts (Secco & Samuroff et al. 2022; Amon et al. 2022); our final data vector has 220 data points whereas the fiducial data vector from DES Y3 has 222 data points.
@@ 4.2 Covariance matrix
The parameter inference described in Section 4.3 below requires estimating the covariance of our measurements. In this work, we generate an analytical covariance using the CosmoCov package (Krause & Eifler 2017; Fang et al. 2020a, b). There are many advantages of using an analytical covariance matrix over one estimated from simulations or from the data itself. A key advantage is that such a covariance matrix does not contain numerical noise and can therefore be robustly inverted to obtain the likelihood. This approach of analytically modeling the covariance is also adopted in the DES Y3 analyses (Friedrich et al. 2021).
CosmoCov uses a two-dimensional FFTLog method to efficiently calculate the real-space, non-Gaussian covariance matrix for galaxy weak lensing (as well as other probes not considered in this work). This covariance can be decomposed as,
@@
where $\mathbf{C}_{\rm G}$ is the standard Gaussian covariance matrix, and the two non-Gaussian contributions include a connected four-point correlation ($\mathbf{C}_{\rm cNG}$) that accounts for the presence of nonlinear structure (Wagner et al. 2015; Barreira & Schmidt 2017a, b) and then a super sample contribution ($\mathbf{C}_{\rm SSC}$) which accounts for cross-correlations between different small-scale modes induced by modes larger than the survey footprint (e.g., Barreira et al. 2018). The non-Gaussian contributions are subdominant for the diagonal elements of $\mathbf{C}$ but are dominant in the off-diagonal terms as their primary effect is in coupling together different modes/scales.
Certain terms in the final covariance can have some sensitivity to the geometry of the survey mask, and some can also have sensitivity to the distribution of galaxies within that mask. We first recast Equation (7) by splitting the covariance into the following terms
@@
where $\mathbf{C}_{\rm SN}$ is the shape-noise term, $\mathbf{C}_{\rm CV/SSC/cNG}$ is the sum of the cosmological terms (i.e., the sole term in the covariance when shape noise is negligible), and $\mathbf{C}_{\rm mix}$ is the mixed term which contains one power each of the cosmological signal and the noise field. In CosmoCov, the terms $\mathbf{C}_{\rm SN}$ and $\mathbf{C}_{\rm SSC}$ (that is, just the super-sample covariance term in $\mathbf{C}_{\rm CV/SSC/cNG}$) include the effects of the survey mask, while the rest ignore it. Friedrich et al. (2021) explicitly show that the impact of the mask on the mixed covariance term, as well as the cosmic variance and connected four-point terms, changes the posterior width by less than $1\%$ (see their Figure 2). The DECADE survey footprint has a similarly regular geometry relative to that of the DES footprint, [^3] The structure of our small-scale mask features is the same as that of DES given we use a similar masking procedure that removes circular regions around bright sources (Paper I). and we therefore follow Friedrich et al. (2021) in assuming the mask can be ignored in the mixed term.
@@
The contribution of the shape noise term is sensitive to the inhomogeneity in the source-galaxy number density. Note that this is not the inhomgeneity in the image quality/depth. Figure 1 illustrates the relevance of this distinction. The left panels show the magnitude limit maps from DECADE and from DES Y3, both generated using the Decasu software. [^4] https://github.com/erykoff/decasu There is a clear visual difference in the homogeneity of these maps. However, this does not directly translate into the middle column, which shows the number density map, $n_{\rm gal}$, of the shape catalog. The catalogs in both DES Y3 and DECADE are defined with a cut $m_{i}<23.5$, which is around the median depth of the DECADE survey. This means the long tail to high magnitude limits (Figure 5 below, or also Figure 1 in Paper I) is alleviated through this cut. Note that this cut was placed in Y3, and is replicated in DECADE, in order to improve the accuracy of our photometric redshift estimates (Myles & Alarcon et al. 2021). The middle column of Figure 1 shows that the distribution of $n_{\rm gal}$ is in fact similar across DECADE and DES Y3. In summary, while the DECADE survey’s observing conditions are more inhomogeneous than DES Y3, the number density of the resulting source-galaxy catalog is similarly homogeneous to that of DES Y3.
@@
Given our survey mask has a regular geometry and similar small-scale feature compared to DES Y3, we rely on the findings on Friedrich et al. (2021) and do not re-validate the $\mathbf{C}_{\rm mix}$ and $\mathbf{C}_{\rm CV/SSC/cNG}$ terms in Equation (7). In this context, we only need to check the impact of inhomogeneity on the shape noise-related terms. As a reminder, the covariance model of Friedrich et al. (2021) included the survey mask in $\mathbf{C}_{\rm SN}$ term, using the method of Troxel et al. (2018), and ignored it in the mixed term (as it contributed only 1% to the posterior scatter, as mentioned above). The variations in galaxy number density within the mask was not included when accounting for the mask. In Appendix C, we confirm that these variations have a minimal impact in the covariance estimate for the DECADE dataset and can therefore be ignored, as was done in DES Y3.
@@
Table 1: Cosmology and nuisance parameters in the baseline $\Lambda$CDM model. Uniform distributions in the range $[a,b]$ are denoted $\mathcal{U}(a,b)$ and Gaussian distributions with mean $\mu$ and standard deviation $\sigma$ are denoted $\mathcal{N}(\mu,\sigma)$. The column “Fiducial” refers to the parameters used for generating the synthetic data vector in all the simulated likelihood tests.
@@ 4.3 Likelihood and sampling
We fit the model above to the cosmic shear measurements of $\xi_{\pm}$ using a Markov Chain Monte Carlo (MCMC) approach. We assume a Gaussian likelihood $L$, with
$\ln L(\xi_{\pm,d}|\boldsymbol{p})=-\frac{1}{2}\bigg{(}\xi_{\pm,d}-\xi_{\pm,m}(\mathbf{p})\bigg{)}\mathbf{C}^{-1}\bigg{(}\xi_{\pm,d}-\xi_{\pm,m}(\mathbf{p})\bigg{)},$ (9)
where $\xi_{\pm,d}$ and $\xi_{\pm,m}$ are the $\xi_{\pm}$ data vector measured in the data and that predicted by the theoretical model of Equation 3, respectively; $\textbf{C}^{-1}$ is the inverse covariance described in Section 4.2; p is a vector of cosmological model parameters and nuisance parameters listed in Table 1. The Bayesian posterior is proportional to the product of the likelihood $L$ and the prior $P$,
$P(\mathbf{p}|\xi_{\pm,d})\propto L(\xi_{\pm,d}|\mathbf{p})P(\mathbf{p}).$ (10)
@@
We perform the parameter inference using the CosmoSIS package (Zuntz et al. 2015). In particular, we use the Nautilus sampler (Lange 2023) for most of our tests, and the Polychord sampler (Handley et al. 2015) for the final cosmological constraints that are shown in Paper IV. The exact parameters we use for the two samplers are listed in Table 2.
@@
Table 2: Parameters used in the different samplers tested in Section 5.4. Nautilus is the primary sampler used in this project. polychord, which is the default sampler in DES, is used to validate our results from Nautilus.
@@ 5 Validation of fiducial model
In this section we validate our model — specifically, testing that it is sufficient for obtaining unbiased constraints — by generating simulated data vectors and analyzing them using the pipeline. We generate a fiducial data vector using the parameters in the “Fiducial” column in Table 1. In particular, the IA parameters have fiducial values taken to be the best-fit parameters from the DES Y3 3$\times$2pt [^5] The $3\times 2$ nomenclature of DES refers to the combination of two-point correlation functions of three probes: galaxy clustering, galaxy galaxy-lensing, and cosmic shear. results (DES Collaboration 2022). The cosmological parameters were first chosen to be generic ($\Omega_{\rm m}=0.3$, $\sigma_{8}=0.8$) in the initial testing, but after unblinding of our final cosmology constrains (see Paper IV) we updated the values to more closely match said constraints. Regardless, the tests in this paper are quite insensitive to the assumed cosmology and to the relatively small shifts we applied pre/post-unblinding.
@@
In Figure 2, we show the resulting cosmological constraint of our fiducial model in the plane spanned by $\Omega_{\rm m}$, $\sigma_{8}$, and $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$. The input cosmology is indicated by the cross-hair and shows our pipeline recovers the input parameters. This is a simple check, but can be non-trivial due to the impact of parameter projection effects (also called “prior volume” effects) in the high-dimensional parameter space as in such cases the marginalized 1D constraints are not guaranteed to be centered at/around the truth. Of the parameters shown in Figure 2, we focus on the derived parameter $S_{8}$, which has been widely adopted in the lensing community since it approximates the most constraining direction in the $\sigma_{8}$-$\Omega_{\rm m}$ plane (Jain & Seljak 1997). Our simulated analysis shows $S_{8}=0.801^{+0.024}_{-0.024}$.
@@
We also overlay the DES Y3 constraints (Secco & Samuroff et al. 2022; Amon et al. 2022) for comparison: $S_{8}=0.759^{+0.025}_{-0.023}$ and $S_{8}=0.772^{+0.018}_{-0.017}$ for the “Fiducial” and ”Optimal” cases, respectively. The “Optimal” case cuts fewer scales from the final analysis. Our focus is the relative size of the contours, and not the relative locations (as we are doing a simulated analysis for the DECADE dataset and have complete freedom in the location of the resulting contour), so we shift the DES posteriors to be centered on the input cosmology of the simulated data vector. The figure shows the DECADE constraints are expected to be similar to the DES Y3 “Fiducial” constraints, and broader that the DES Y3 “Optimal” constraints. Our scale cuts match the former setup. We also note that all DES constraints shown here include information from shear ratios (Sánchez & Prat et al. 2022). We perform a more detailed comparison of the DECADE and DES Y3 constraints in Paper IV and find the DECADE result slightly outperform the DES Y3 “Fiducial” result once all analysis choices are matched.
@@
Figure 2: Forecasted constraints on $\Omega_{\rm m}$, $\sigma_{8}$ and $S_{8}$ using a simulated data vector corresponding to the DECADE dataset. The cross-hair indicates the input cosmology, showing that our pipeline adequately recovers the input values. The DES Y3 constraints from Secco & Samuroff et al. (2022) and Amon et al. (2022) are overlaid for comparison. To enhance comparisons of constraining power, we shift the DES posteriors to be centered on the simulated data vector’s input cosmology. Both the fiducial and the optimal constraints from DES include information from shear ratios (Sánchez & Prat et al. 2022), resulting in tighter constraining power and slightly different degeneracy directions.
@@
In the following subsections we perform a number of stress-tests on the model. We summarize all the tests in Figure 3, where we show the marginal constraints in $S_{8}$, $\Omega_{\rm m}$ and $\sigma_{8}$, and compare it with the fiducial constraints in Figure 2. First, Section 5.1 confirms that our fiducial scale cuts (derived in Section 4.1) alleviate biases from baryonic imprints in the data vector. Next, Section 5.2 discusses shifts in constraints when assuming different models for the nonlinear matter power spectrum and the IA. In Section 5.3 we examine the sensitivity of our constraints to the priors on the nuisance parameters, and in Section 5.4, the sensitivity of the constraints to the choice of sampling methods.

### v2

@@ 3.3 Shear and redshift calibration uncertainties
The two main sources of systematic uncertainties on our measurements are those associated with estimates of the shear and of the redshift distribution. We follow the implementation from DES Y3 (Secco & Samuroff et al. 2022; Amon et al. 2022) — which is consistent with those of numerous other works (e.g., Asgari et al. 2021; Dalal et al. 2023) — in the modeling of these uncertainties. We now detail our implementation below.
For any shear measurement algorithm, a calibration step is needed for the measured shear. This calibration, which is obtained either from simulations or the data itself, carries an associated uncertainty. In the weak lensing regime, the residual impact of any biases on the measured shear, $\gamma_{\textrm{obs}}$, is modeled with both multiplicative $m_{i}$ and additive components $c_{i}$ (Huterer et al. 2006; Heymans et al. 2006; Bridle & King 2007), that scale the true shear $\gamma_{\textrm{true}}$ as
@@
where $m_{i}$ and $c_{i}$ are constants corresponding to the $i^{\rm th}$ redshift bin. The additive, residual bias often has a sufficiently small amplitude to not be relevant (e.g., MacCrann & Becker et al. 2022; Amon et al. 2022). Furthemore, the shear measurements of our catalog are mean-subtracted (Paper I) following the approach in other works (e.g., Zuntz & Sheldon et al. 2018; Gatti & Sheldon et al. 2021; Yamamoto & Becker et al. 2025). This naturally corrects for such additive biases in the data. The multiplicative residual bias, however, scales the amplitude of $\xi_{\rm\pm}$ and thus propagates directly to key cosmological constraints. As a result, accurate estimates of $m_{i}$ are a necessity. We have made such estimates for our shear catalog using a suite of simulated images of the DECADE survey, as described in detail in Paper I. The simulations inform priors for $m_{i}$ which are listed in Table 1. We note that these simulations are data-driven and reproduce the range of observing conditions seen in the real data; see Paper I.
For the redshift distribution, $n(z)$, we also follow DES Y3 in using a shift in the mean redshift, $\Delta z_{i}$, per bin that then shifts the measured $n(z)$ for each bin $i$ as
$n_{i}(z)=n_{{\rm orig},i}\left(z-\Delta z_{i}\right).$ (8)
where $n_{{\rm orig},i}$ is the original distribution. This parametrization keeps the shape of the $n(z)$ unchanged. While one could explore more complicated parametrizations of the $n(z)$ uncertainties, the above model is sufficient for the statistical power of our data (which is similar to DES Y3). In Paper II, we derived the priors on $\Delta z_{i}$ using the self-organizing map photometric redshifts (SOMPZ) method while accounting for a variety of uncertainty contributions including cosmic variance, shot noise, redshift bias, and photometric zeropoint uncertainties. In the Appendix A of Paper II, we also derived estimates of the $n(z)$ from combining SOMPZ with an independent estimate, clustering redshifts (WZ). The latter method includes information from the spatial clustering between our shear sample and spectroscopic sources. The former, SOMPZ-only method is our fiducial choice, while we use the latter SOMPZ plus WZ method as a cross-check. The calibrated priors on $\Delta z_{i}$ from the fiducial setup are listed in Table 1. Our redshift distributions generally span $0<z<1.5$. See Figure 9 of Paper II for more details.
@@ 4.1 Scale cuts
As we discussed in Section 3.1, baryons can influence the matter power spectrum on a wide range of scales, presenting a significant source of uncertainty. For example, feedback processes from AGN and supernovae can heat environments and suppress matter clustering, while metal enrichment may offer cooling channels that increase small-scale clustering (e.g., Chisari et al. 2018; Schneider et al. 2019). To reduce the impact of these baryonic imprints, we follow Secco & Samuroff et al. (2022) and Amon et al. (2022) in removing shear measurements whose signal can have a significant contribution from baryon evolution. In particular, we use a simulated data vector which is generated using the matter-power spectrum measured in the “AGN” model of the OverWhelmingly Large Simulations (OWLS, Schaye et al. 2010, see their Table 2). The current estimates of the power-spectrum suppression from data are statistically consistent with this OWLS model (e.g., Bigwood et al. 2024, see their Figure 1 and 4).
[…]
@@
  • Of the 20 selected points from the previous step, we find the data point with the highest $\Delta\chi^{2}$ and remove it. We repeat this iteratively until the $\Delta\chi^{2}$ between the contaminated and uncontaminated data vectors, after masking/scale cuts, is $\leq 0.3$. This $\Delta\chi^{2}$ between the two data vectors is determined using the entire covariance matrix (and not just its diagonal component).
The $\xi_{-}$ shear measurements are more heavily affected by baryon contamination and thus require more impactful scale cuts. In practice, we remove a similar number of points as the fiducial DES Y3 scale cuts (Secco & Samuroff et al. 2022; Amon et al. 2022); our final data vector has 220 data points whereas the fiducial data vector from DES Y3 has 222 data points. The unmasked datavector has 400 points — 20 angular separations per tomographic bin and per $\xi_{+}$ and $\xi_{-}$. Our data vector, and the associated scale cuts per tomographic bin, is shown in Figure 3 of Paper IV.
@@ 4.2 Covariance matrix
The parameter inference described in Section 4.3 below requires estimating the covariance of our measurements. We now describe our approach to generating this covariance, and also check that the fiducial framework is sufficient for inhomogeneous datasets of the kind considered in this work.
In this work, we generate an analytical covariance using the CosmoCov package (Krause & Eifler 2017; Fang et al. 2020a, b). See references therein for the analytic formalism and validations of this formalism. There are many benefits to using an analytical covariance matrix over one estimated from simulations or from data resampling (e.g., jackknife covariances). A key advantage is that such a covariance matrix does not contain numerical noise and can therefore be robustly inverted to obtain the likelihood. This approach of analytically modeling the covariance is also adopted in the DES Y3 analyses (Friedrich et al. 2021).
CosmoCov uses a two-dimensional FFTLog method (Hamilton 2000) to efficiently calculate the real-space, non-Gaussian covariance matrix for galaxy weak lensing (as well as other probes not considered in this work). This covariance can be decomposed as,
@@
where $\mathbf{C}_{\rm G}$ is the standard Gaussian covariance matrix, and the two non-Gaussian contributions include a connected four-point, non-Gaussian correlation ($\mathbf{C}_{\rm cNG}$) that accounts for the presence of nonlinear structure (Wagner et al. 2015; Barreira & Schmidt 2017a, b) and then a super sample contribution ($\mathbf{C}_{\rm SSC}$) which accounts for cross-correlations between different small-scale modes induced by modes larger than the survey footprint (e.g., Barreira et al. 2018). The non-Gaussian contributions are subdominant for the diagonal elements of $\mathbf{C}$ but are dominant in the off-diagonal terms as their primary effect is in coupling together different modes/scales.
Certain terms in the final covariance can have some sensitivity to the geometry of the survey mask, and some can also have sensitivity to the distribution of galaxies within that mask (we detail the exact terms further below). We first recast Equation (9) — to distinguish the contributions from noise and those from cosmic variance (CV) — by splitting the covariance into the following terms
@@
where $\mathbf{C}_{\rm SN}$ is the shape-noise term, $\mathbf{C}_{\rm CV/SSC/cNG}$ is the sum of the cosmological terms (i.e., the sole term in the covariance when shape noise is negligible), and $\mathbf{C}_{\rm mix}$ is the mixed term which contains one power each of the cosmological signal and the noise field. In CosmoCov, the terms $\mathbf{C}_{\rm SN}$ and $\mathbf{C}_{\rm SSC}$ (that is, just the super-sample covariance term in $\mathbf{C}_{\rm CV/SSC/cNG}$) include the effects of the survey mask, while the $\mathbf{C}_{\rm CV}$ and $\mathbf{C}_{\rm cNG}$ terms ignore it. Friedrich et al. (2021) explicitly show that the impact of the mask on the mixed covariance term, as well as the cosmic variance and connected four-point terms, changes the posterior width by less than $1\%$ (see their Figure 2). The DECADE survey footprint has a similarly regular geometry relative to that of the DES footprint, [^4] The structure of our small-scale mask features is the same as that of DES given we use a similar masking procedure that removes circular regions around bright sources (Paper I). and we therefore follow Friedrich et al. (2021) in assuming the mask can be ignored in the mixed term.
@@
The contribution of the shape noise term is sensitive to the inhomogeneity in the source-galaxy number density. Note that this is not the inhomogeneity in the image quality/depth. Figure 1 illustrates the relevance of this distinction. The left panels show the magnitude limit maps from DECADE and from DES Y3, both generated using the Decasu software. [^5] https://github.com/erykoff/decasu There is a clear visual difference in the homogeneity of these maps. However, this does not directly translate into the middle column, which shows the number density map, $n_{\rm gal}$, of the shape catalog. The catalogs in both DES Y3 and DECADE are defined with a cut on the $i$-band magnitude, $m_{i}<23.5$, which is around the median depth of the DECADE survey. This means the long tail to high magnitude limits (Figure 5 below, or also Figure 1 in Paper I) is alleviated through this cut. Note that this cut was placed in DES Y3, and is replicated in DECADE, in order to improve the accuracy of our photometric redshift estimates (Myles & Alarcon et al. 2021). The middle column of Figure 1 shows that the variation in $n_{\rm gal}$ is in fact similar across DECADE and DES Y3. In summary, while the DECADE survey’s observing conditions are more inhomogeneous than DES Y3, the number density of the resulting source-galaxy catalog is similarly homogeneous to that of DES Y3.
@@
Given our survey mask has a regular geometry and similar small-scale feature compared to DES Y3, we rely on the findings on Friedrich et al. (2021) and do not revalidate the $\mathbf{C}_{\rm mix}$ and $\mathbf{C}_{\rm CV/SSC/cNG}$ terms in Equation (9). In this context, we only need to check the impact of inhomogeneity on the shape noise-related terms. As a reminder, the covariance model of Friedrich et al. (2021) included the survey mask in $\mathbf{C}_{\rm SN}$ term, using the method of Troxel et al. (2018), and ignored it in the mixed term (as it contributed only 1% to the posterior scatter, as mentioned above). The variations in galaxy number density within the mask was not included when accounting for the mask. In Appendix C, we confirm that these variations have a minimal impact in the covariance estimate for the DECADE dataset and can therefore be ignored, as was done in DES Y3.
@@
[…]
@@ 4.3 Likelihood and sampling
We fit the model above to the cosmic shear measurements of $\xi_{\pm}$ using a Markov Chain Monte Carlo (MCMC) approach. We assume a Gaussian likelihood $\mathcal{L}$, with
$\ln\mathcal{L}(\xi_{\pm,d}|\boldsymbol{p})\propto-\frac{1}{2}\bigg(\xi_{\pm,d}-\xi_{\pm,m}(\mathbf{p})\bigg)\mathbf{C}^{-1}\bigg(\xi_{\pm,d}-\xi_{\pm,m}(\mathbf{p})\bigg),$ (11)
where $\xi_{\pm,d}$ and $\xi_{\pm,m}$ are the $\xi_{\pm}$ data vector measured in the data and that predicted by the theoretical model of Equation 3, respectively; $\textbf{C}^{-1}$ is the inverse covariance described in Section 4.2; p is a vector of cosmological model parameters and nuisance parameters listed in Table 1. The Bayesian posterior is proportional to the product of the likelihood $\mathcal{L}$ and the prior $P$,
$P(\mathbf{p}|\xi_{\pm,d})\propto\mathcal{L}(\xi_{\pm,d}|\mathbf{p})P(\mathbf{p}).$ (12)
@@
We perform the parameter inference using the CosmoSIS package (Zuntz et al. 2015). In particular, we use the Nautilus sampler (Lange 2023) for most of our tests, and have cross-checked that the results are consistent with those obtained using the Polychord sampler (Handley et al. 2015). We also produce results using the MultiNest sampler (Feroz et al. 2009) for an additional comparison point. The exact parameters we use for the samplers are listed in Table 2.
@@
Table 2: Parameters used in the different samplers tested in Section 5.4. Nautilus is the primary sampler used in this project. Polychord, which is the default sampler in DES analyses, is used to validate our results obtained with Nautilus. We also use Multinest as an additional comparison point.
@@ 5 Validation of fiducial model
[…]
@@
In Figure 2, we show the resulting cosmological constraint of our fiducial model in the plane spanned by $\Omega_{\rm m}$, $\sigma_{8}$, and $S_{8}\equiv\sigma_{8}\sqrt{\Omega_{\rm m}/0.3}$. The input cosmology is indicated by the cross-hair and shows our pipeline recovers the input parameters. This is a simple check, but can be non-trivial due to the impact of parameter projection effects (also called “prior volume” effects) — caused by a high-dimensional parameter space and/or sufficiently asymmetric posteriors — as in such cases the marginalized 1D constraints are not guaranteed to be centered at/around the truth. Of the parameters shown in Figure 2, we focus on the derived parameter $S_{8}$, which has been widely adopted in the lensing community since it approximates the most constraining direction in the $\sigma_{8}$-$\Omega_{\rm m}$ plane (Jain & Seljak 1997). Our simulated analysis shows $S_{8}=0.801^{+0.024}_{-0.024}$, where the uncertainties denote the 68% confidence interval computed from the posterior.
@@
We also overlay the DES Y3 constraints (Secco & Samuroff et al. 2022; Amon et al. 2022) for comparison: $S_{8}=0.759^{+0.025}_{-0.023}$ and $S_{8}=0.772^{+0.018}_{-0.017}$ for the “Fiducial” and ”Optimal” cases, respectively. The “Optimal” case cuts fewer scales from the final analysis, and therefore provides better constraints. Our focus is the relative size of the contours, and not the relative locations (as we are doing a simulated analysis for the DECADE dataset and have complete freedom in the location of the resulting contour), so we shift the DES posteriors to be centered on the input cosmology of the simulated data vector. The figure shows the DECADE constraints are expected to be similar to the DES Y3 “Fiducial” constraints, and broader that the DES Y3 “Optimal” constraints. Our scale cuts match the former setup. We also note that all DES constraints shown here include information from shear ratios (Sánchez & Prat et al. 2022). We perform a more detailed comparison of the DECADE and DES Y3 constraints in Paper IV and find the DECADE result slightly outperform the DES Y3 “Fiducial” result once all analysis choices are matched.
@@
Figure 2: Forecasted constraints on $\Omega_{\rm m}$, $\sigma_{8}$ and $S_{8}$ using a simulated data vector corresponding to the DECADE dataset (red). The cross-hair indicates the input cosmology, showing that our pipeline adequately recovers the input values. The DES Y3 constraints from Secco & Samuroff et al. (2022) and Amon et al. (2022) are overlaid for comparison. To enhance comparisons of constraining power, we shift the DES posteriors to be centered on the simulated data vector’s input cosmology. Both the fiducial and the optimal constraints from DES include information from shear ratios (Sánchez & Prat et al. 2022), resulting in tighter constraining power and slightly different degeneracy directions.
@@
In the following subsections we perform a number of stress-tests on the model. We summarize all the tests in Figure 3, where each line shows the marginal $1\sigma$ and $2\sigma$ intervals in $S_{8}$, $\Omega_{\rm m}$ and $\sigma_{8}$ for a given modeling scenario, and compare it with the intervals of the fiducial constraints (vertical shaded bands) that was also shown in Figure 2. First, Section 5.1 confirms that our fiducial scale cuts (derived in Section 4.1) alleviate biases from baryonic imprints in the data vector. Next, Section 5.2 discusses shifts in constraints when assuming different models for the nonlinear matter power spectrum and the IA. In Section 5.3 we examine the sensitivity of our constraints to the priors on the nuisance parameters, and in Section 5.4, the sensitivity of the constraints to the choice of sampling methods.